Consider a value to be significantly low if its z score less than or equal to minus2 or consider a value to be significantly high if its z score is greater than or equal to 2. A test is used to assess readiness for college. In a recent​ year, the mean test score was 20.2 and the standard deviation was 5.2. Identify the test scores that are significantly low or significantly high. What test scores are significantly​ low? Select the correct answer below and fill in the answer​ box(es) to complete your choice.
Answer #1

Answer

It is given in the question that values which are less than 2 z scores or more than 2 z scores are significantly low values and high values respectively

This means that we have to find values that are two standard deviations away from mean

given that

mean = 20.2 and standard deviation = 5.2

Significantly low value = mean - 2*sd

= 20.2 - 2*5.2

= 9.8 or below

and

significantly high value = mean + 2*sd

= 20.2 + 2*5.2

= 30.6 or above